I need help graphing this problem
Graph the line
3x+3y+11=0

Answer by pug912(10)

to graph the line 3x+3y+11=0
put x=0, then y = -11/3
put y =0, then x = -11/3
so the line should pass through these two points (0,-11/3) and (-11/3,0).
to draw the line, just mark two points (0,-11/3) and (-11/3,0) in your graph paper and connect these two points using line.